# Datalog-Expressibility for Monadic and Guarded Second-Order Logic

### v2

@@ 6 Nested Guarded Queries
Frontier-guarded Datalog is a fragment of Datalog that is contained in GSO. Guarded queries (GQ) and the more expressive Nested guarded queries (GQ+) are extensions of frontier-guarded Datalog that have been introduced by Bourhis, Krötzsch, and Rudolph [18]; the definitions will be recalled below. GQ is also strictly more expressive than Nested Monadically Defined Queries (Nemodeq), Monadically defined queries (Modeq) [42], and the recently introduced almost monadic queries [43]. Both GQ and GQ+ are contained in the intersection of Datalog and GSO. We show that GQ+ does not contain all queries from the intersection of Datalog and MSO (and hence, neither does GQ, frontier-guarded Datalog, Nemodeq, Modeq, and almost monadic Datalog): there are CSPs that are both in MSO and in Datalog, but not in GQ+ (Corollary 6.51).
@@ 6.1 Frontier-Guarded Datalog
A rule of a Datalog program is called frontier-guarded if all variables of the head appear in a single EDB atom in the rule body. [^4] The term frontier-guarded was originally introduced for existential rules (also referred to as tuple-generating dependencies or Datalog±), an expressive extension of Datalog. Frontier-guarded Datalog is then simply the syntactic intersection of frontier-guarded existential rules and Datalog. The name guarded is reserved for rules where all variables of the rule appear in a single EDB atom in the rule body. A Datalog program is called frontier-guarded if all its rules are frontier-guarded. Note that every monadic Datalog program is frontier-guarded. Also the Datalog program from the proof of Proposition 2.5 is guarded.
@@
  • additionally adding the conjunct $\neg{\sf goal}$ to $\Phi$.
@@
We will now show that, if $\Pi$ is frontier-guarded, then $\Phi_{\Pi}$ can be expressed by a GSO sentence. To this end, we establish that $\Phi_{\Pi}$ has the same meaning under standard and under guarded semantics, from which the desired result follows via Proposition 2.4. Toward a contradiction, assume some structure $\mathfrak{A}$ satisfies $\Phi_{\Pi}$ under standard but not under guarded semantics (as $\Phi_{\Pi}$ is preceded only by existentially quantified relation symbols, this is the only possibility). Assume that instantiating the IDBs with the relations $R_{1},\ldots,R_{k}$ over $A$ witnesses that $\mathfrak{A}$ satisfies $\Phi_{\Pi}$ under standard semantics. Obtain now the guarded relations $R^{\prime}_{1},\ldots,R^{\prime}_{k}$ by removing all unguarded tuples from $R_{1},\ldots,R_{k}$, respectively. The desired contradiction is now obtained by arguing that instantiating the IDBs with the relations $R^{\prime}_{1},\ldots,R^{\prime}_{k}$ witnesses that $\mathfrak{A}$ satisfies $\Phi_{\Pi}$ under guarded semantics: First, we note that $\neg{\sf goal}$ is still satisfied. Second, for any rule $\psi\;{:}{-}\;\phi_{1},\dots,\phi_{m}$ of $\Pi$, fixing any variable assignment of $\bar{x}$ with elements of $A$, we obtain that the truth of $\psi\vee\neg\phi_{1}\vee\cdots\vee\neg\phi_{m}$ under the instantiation with $R_{1},\ldots,R_{k}$ implies its truth under the instantiation with $R^{\prime}_{1},\ldots,R^{\prime}_{k}$: On the one hand, should any $\neg\phi_{i}$ be true, this is immediate (both for EDB and IDB atoms). On the other hand, in case all $\neg\phi_{i}$ are false, then so is the one $\neg\phi_{j}$ where $\phi_{j}=\phi_{j}(\bar{y})$ is the rule’s “frontier guard”, i.e., it consists of an EDB atom where $\bar{y}$ contains at least all the variables of $\bar{z}$ from $\psi(\bar{z})=\psi$. Truth of $\phi_{j}(\bar{y})$ under the chosen assignment implies that the tuple assigned to $\bar{y}$ is guarded and therefore also the tuple assigned to $\bar{z}$ must be guarded. Yet, then, truth of $\psi(\bar{z})$ under the two IDB instantiations coincides, hence $\psi(\bar{z})$ must hold.
However, not every problem in GSO which is in Datalog can be expressed by a frontier-guarded Datalog program. To prove this we need the following definition.
@@
Let $\mathfrak{B}$ be an $\omega$-categorical structure with finite relational signature $\tau$ and $k\in{\mathbb{N}}$. Let $s$ be the maximal arity of $\tau$. Then the guarded canonical Datalog program of width $k$ for the complement of $\operatorname{CSP}(\mathfrak{B})$ is the subset of the canonical Datalog program of width $(s,k)$ for the complement of $\operatorname{CSP}(\mathfrak{B})$ which contains all rules that are guarded.
If $\Pi$ the the guarded canonical Datalog program of width $k$ for the complement of $\operatorname{CSP}(\mathfrak{B})$ then it follows from Proposition 6.5 below that $\llbracket\Pi^{\prime}\rrbracket\subseteq\llbracket\Pi\rrbracket$ for every frontier-guarded Datalog program $\Pi^{\prime}$ of width $(k,k)$ which is sound for the complement of $\operatorname{CSP}(\mathfrak{B})$. We need to adapt the existential pebble game on $\mathfrak{A}$ and $\mathfrak{B}$ to the guarded setting as well. In our informal definition of the existential pebble game, we only change the rules for Spoiler: we additionally require that when Spoiler removes pebbles from $\mathfrak{A}$, then all the pebbles that remain on $\mathfrak{A}$ must be guarded, i.e., there must exist a tuple in a relation of $\mathfrak{A}$ such that these pebbles must be from the entries of this tuple. The resulting game will be called the existential guarded $k$ pebble game. Formally, we again work exclusively with the concept of a winning strategy ${\mathcal{H}}$ for Duplicator (similarly as in Definition 4.11). A partial homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$ is called guarded if there exists a tuple in a relation of $\mathfrak{A}$ whose entries contain $\operatorname{dom}(h)$.
@@
The following can be shown analogously to the proof of Theorem 4 in [12]. We omit the proof because all the ideas are already present in the proof of a more interesting theorem that we present in full detail in Section 6.4 (Theorem 6.26).
@@
  • The canonical frontier-guarded Datalog program of width $k$ for the complement of $\operatorname{CSP}(\mathfrak{B})$ does not derive ${\sf goal}$ on $\mathfrak{A}$.
@@
Proof 6.7.
Let $k\in{\mathbb{N}}$. By Proposition 6.5, it suffices to show that there exists a structure $\mathfrak{A}$ which has no homomorphism to $({\mathbb{Q}};<)$ but Duplicator has a winning strategy for the existential guarded $k$ pebble game on $\mathfrak{A},({\mathbb{Q}};<)$. Let $\mathfrak{A}$ be a directed cycle of length at least $k+1$. Let ${\mathcal{H}}$ be the set of all partial homomorphisms $h$ from $\mathfrak{A}$ to $({\mathbb{Q}};<)$ with a domain of size at most $k$. We claim that $\mathcal{H}$ is a winning strategy for the existential guarded $(\ell,k)$-pebble game. Clearly, ${\mathcal{H}}$ is closed under restrictions. Now, let $h$ be a partial homomorphism whose domain is contained in a tuple from a relation of $\mathfrak{A}$, that is, contained in $\{a_{1},a_{2}\}$ for two subsequent elements of $A$ on the cycle. Let $a_{3},\dots,a_{k}$ be elements of $A$. Since $|A|\geq k+1$ there must be some $b\in A\setminus\{a_{1},\dots,a_{k}\}$. Let $\prec$ be the order on $a_{1},\dots,a_{k}$ in which these elements are traversed on the cycle, starting with $b$. Note that $\prec$ is a linear extension of $<^{\mathfrak{A}}$. We may then extend $h$ to a partial homomorphism $h^{\prime}$ from $\mathfrak{A}$ to $({\mathbb{Q}};<)$ such that $h(a_{i})<h(a_{j})$ if $a_{i}\prec a_{j}$, for all $i,j\in\{1,\dots,k\}$. Note that $h^{\prime}\in{\mathcal{H}}$; this finishes the proof.
@@ 6.2 Flag and Check
A general mechanism, called flag and check, to turn a query language into a potentially more powerful query language has been introduced by Bourhis, Krötzsch, and Rudolph [18], generalising earlier work of Rudolph and Krötzsch [42]. The idea is best described for some fundamental constraint satisfaction problems. It is easy to see that there is no $k$ such that $\operatorname{CSP}(\{0,1\};\neq)$ (i.e., graph 2-colorability) is solved by a Datalog program of width $(1,k)$. However, to decide whether a graph is 2-colorable, it suffices to show that there is no vertex $v$ which can reach $v$ via a path of odd length. Note that such a ‘check’ can be performed by a Datalog program of width $(1,2)$ once a ‘flag’ has been put on $v$. Similarly, there is no $k\in{\mathbb{N}}$ such that $\operatorname{CSP}({\mathbb{Q}};<)$ (i.e., digraph acyclicity) can be solved by a Datalog program of width $(1,k)$ [7]. However, to decide whether a graph contains a directed cycle, it suffices to find a vertex $v$ which can reach $v$ by a directed path with at least one edge. Again, after $v$ has been found and ‘flagged’ such a computation can be performed by a Datalog program of width $(1,2)$.
@@
Let $\tau$ be a finite set of relation and constant symbols. A flag-and-check $\tau$-program ($\tau$-FCP) of arity $m$ is a set of Datalog rules $\Pi$ with EDBs $\tau\cup\{\lambda_{1},\dots,\lambda_{m}\}$, where $\lambda_{1},\dots,\lambda_{m}$ are new constant symbols, and the IDBs $\{{\sf goal},P_{1},\dots,P_{k}\}$ where ${\sf goal}$ is a distinguished predicate of arity 00. If $\mathfrak{A}$ is a $\tau$-structure, then $\Pi^{\mathfrak{A}}$ is the set of all tuples $(a_{1},\dots,a_{m})\in A^{m}$ such that $\Pi$ derives ${\sf goal}$ on the $\tau\cup\{\lambda_{1},\dots,\lambda_{m}\}$-expansion $\mathfrak{A}^{\prime}$ of $\mathfrak{A}$ where $\lambda^{\mathfrak{A}^{\prime}}_{i}=a_{i}$ for all $i\in\{1,\dots,m\}$.
@@
We omit the reference to $\tau$ in the notation if the reference is clear from the context. It has been shown that every problem in Modeq can be expressed in MSO [42]; similarly, we obtain the following.
@@
An example of a query that can be expressed in GQ, but not in Modeq can be found in [18] (Example 2). Certain flag-and-check programs have been studied in the context of the complexity of constraint satisfaction under the name peek arc consistency, extending the famous (hyper-) arc consistency procedure [9]. In the following, $\tau$ denotes a finite relational signature.
Definition 6.11.
For example, the PAC procedure for $\operatorname{CSP}(\{0,1\};\neq)$ derives ${\sf goal}$ on a finite structure $\mathfrak{A}$ if and only if $\mathfrak{A}$ has no homomomorphism to ($(\{0,1\};\neq)$; similarly, the PAC procedure solves $\operatorname{CSP}({\mathbb{Q}};<)$. We prove that the PAC procedure can be expressed as a Modeq.
Lemma 6.12.
Proof 6.13.
Theorem 3.1 implies that we may assume without loss of generality that the automorphism group of $\mathfrak{B}$ has finitely many orbits $O_{1},\dots,O_{n}$; choose representatives $b_{1},\dots,b_{n}$ from these orbits. For every $i\leq n$, let $\Pi_{i}$ be the Datalog program for the hyperarc consistency procedure for $\operatorname{CSP}(\mathfrak{B},\{b_{i}\})$ (which has width $(1,p)$; see Section 8.4 in [8]). Let $C_{i}$ be the unary IDB of $\Pi_{i}$ for the relation $\{b_{i}\}$, and let $\Pi^{\prime}_{i}$ be the Datalog program obtained from $\Pi_{i}$ by adding the rule $C_{i}(\lambda_{1})\;{:}{-}$ (without precondition). Since expressibility by Datalog programs of width $(1,p)$ is closed under intersection (Lemma 4.3), there is a Datalog program $\Pi$ of width $(1,p)$ expressing $\Pi^{\prime}_{1}\wedge\cdots\wedge\Pi^{\prime}_{n}$. Also note that $\Theta:=\exists x.\Pi$ holds on a $\tau$-structure $\mathfrak{A}$ if and only if there exists an element $a\in A$ such that for every $i\leq n$ the program $\Pi^{\prime}_{i}$ derives ${\sf goal}$ on $(\mathfrak{A},\{a\})$, which by definition is the case if and only if the PAC procedure derives ${\sf goal}$  on $\mathfrak{A}$. This shows the statement.
Example 6.14.
The following is based on an example from [22] of a structure such that the complement of the CSP of this structure cannot be solved by PAC, but can be solved by singleton arc consistency (SAC), and hence is in Datalog. We will show below that it is even in Modeq. Let $\mathfrak{B}$ be the structure with domain $\{-1,0,1,2\}$ and relations
$R^{\mathfrak{B}}$$\coloneqq\{-1,0,1,2\}^{2}\setminus\{(-1,-1)\}$
$S^{\mathfrak{B}}$$\coloneqq\{(0,1),(1,2),(2,0),(-1,-1)\}$
$R^{\mathfrak{A}}$$\coloneqq\{(a,a^{\prime})\}$
Then there is no homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$, but the PAC procedure for $\operatorname{CSP}(\mathfrak{B})$ does not derive ${\sf goal}$ on $\mathfrak{A}$ [22]. However, the problem can be solved by the Modeq $\exists x_{1},x_{2}.\,\Pi$ where $\Pi$ is the following monadic flag-and-check program of arity two.
[…]
$U_{1}(z)$ $\;{:}{-}\;U_{0}(y),S(y,z)$ $U_{2}(y)$ $\;{:}{-}\;U_{0}(z),S(y,z)$
$U_{2}(z)$ $\;{:}{-}\;U_{1}(y),S(y,z)$ $U_{0}(y)$ $\;{:}{-}\;U_{1}(z),S(y,z)$
$U_{0}(z)$ $\;{:}{-}\;U_{2}(y),S(y,z)$ $U_{1}(y)$ $\;{:}{-}\;U_{2}(z),S(y,z)$
${\sf goal}$ $\;{:}{-}\;U_{0}(y),U_{1}(y),R(y,z),U_{0}(z),U_{1}(z)$
Claim. $\exists x_{1},x_{2}.\,\Pi$ evaluates to true on a finite $\{R,S\}$-structure $\mathfrak{A}$ if and only if there is no homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$. First suppose that there are $a_{1},a_{2}\in A$ such that $\Pi$ derives ${\sf goal}$ on the $\{\lambda_{1},\lambda_{2}\}$-expansion $\mathfrak{A}^{\prime}$ of $\mathfrak{A}$ where $\lambda_{i}^{\mathfrak{A}_{1}}=a_{i}$ for $i\in\{1,2\}$. This means that there are $b_{1},b_{2}\in A$ such that $(b_{1},b_{2})\in R^{\mathfrak{A}}$, and there are paths of net length 0 and 1 from $a_{1}$ to $b_{1}$, and paths of net length 0 and 1 from $a_{2}$ to $b_{2}$. In general, the existence of paths of net length 00 and $1$ between two vertices implies that both have to be mapped to $-1$. Therefore, any homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$ must map both $b_{1}$ and $b_{2}$ to $-1$, which is impossible since $(b_{1},b_{2})\in R^{\mathfrak{A}}$.
Now suppose that for any $a_{1},a_{2}\in A$ the program $\Pi$ does not derive ${\sf goal}$  on the structure $\mathfrak{A}^{\prime}$ defined above. That is, for every $(b_{1},b_{2})\in R^{\mathfrak{A}}$ there exists $i\in\{1,2\}$ such that the for all $u\in A$ that are connected to $b_{i}$ in the graph defined by $S^{\mathfrak{A}}$ there is no path of net length 00 and no path of net length $1$ to $b_{i}$. Hence, there must exist a path of net length $2$ from $u$ to $b_{i}$. We then define $h(u):=b_{i}+2\mod 3$. For all other $u\in A$, we define $h(u):=-1$. It is straightforward to verify that $h$ is a homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$. $\triangle$
An example of a Datalog query in MSO which cannot be expressed in Modeq has been presented in [42] (Example 6); in fact, the same query is not even expressible in GQ, as we will show in Section 6.4 (Example 6.30), so we present it in more detail in the following.
@@
$\exists U_{2},U_{3}\exists z\forall u,v\big{(}(C(u,z)$ $\Rightarrow U_{3}(u))$

### v3

@@ 5 Nested Guarded Queries
In this section we revisit logics that express problems that are contained in both GSO and Datalog, and then prove that all of them fail to express all problems that lie in the intersection of GSO and Datalog. Frontier-guarded Datalog is a fragment of Datalog that is contained in GSO. Guarded queries (GQ) and the more expressive Nested guarded queries (GQ+) are extensions of frontier-guarded Datalog that have been introduced by Bourhis, Krötzsch, and Rudolph [21]; the definitions will be recalled below. GQ is also strictly more expressive than Nested Monadically Defined Queries (Nemodeq), Monadically defined queries (Modeq) [43], and the recently introduced almost monadic queries [44]. Both GQ and GQ+ are contained in the intersection of Datalog and GSO. We show that GQ+ does not contain all queries from the intersection of Datalog and MSO (and hence, neither do GQ, frontier-guarded Datalog, Nemodeq, Modeq, and almost monadic Datalog): there are CSPs whose complement is both in MSO and in Datalog, but not in GQ+ (Corollary 5.39).
@@ 5.1 Frontier-Guarded Datalog
A rule of a Datalog program is called frontier-guarded if all variables of the head appear in a single EDB atom in the rule body. [^6] The term frontier-guarded was originally introduced for existential rules (also referred to as tuple-generating dependencies or Datalog±), an expressive extension of Datalog. Frontier-guarded Datalog is then simply the syntactic intersection of frontier-guarded existential rules and Datalog. The name guarded is reserved for rules where all variables of the rule (and not just the variables of the head) appear in a single EDB atom in the rule body. A Datalog program is called frontier-guarded if all its rules are frontier-guarded. Note that every monadic Datalog program is frontier-guarded. Also the Datalog program from the proof of Proposition 2.5 is frontier-guarded.
@@
  • additionally adding the conjunct $\neg{\sf goal}$ to $\Phi_{\Pi}$.
@@
We will now show that, if $\Pi$ is frontier-guarded, then $\Phi_{\Pi}$ can be expressed by a GSO sentence. To this end, we establish that $\Phi_{\Pi}$ has the same meaning under standard and under guarded semantics, from which the desired result follows via Proposition˜2.4. Toward a contradiction, assume some structure $\mathfrak{A}$ satisfies $\Phi_{\Pi}$ under standard but not under guarded semantics (as $\Phi_{\Pi}$ is preceded only by existentially quantified relation symbols, this is the only possibility). Assume that instantiating the IDBs with the relations $R_{1},\ldots,R_{k}$ over $A$ witnesses that $\mathfrak{A}$ satisfies $\Phi_{\Pi}$ under standard semantics. Obtain now the guarded relations $R^{\prime}_{1},\ldots,R^{\prime}_{k}$ by removing all unguarded tuples from $R_{1},\ldots,R_{k}$, respectively. The desired contradiction is now obtained by arguing that instantiating the IDBs with the relations $R^{\prime}_{1},\ldots,R^{\prime}_{k}$ witnesses that $\mathfrak{A}$ satisfies $\Phi_{\Pi}$ under guarded semantics: First, we note that $\neg{\sf goal}$ is still satisfied. Second, for any rule $\psi\;{:}{-}\;\phi_{1},\dots,\phi_{m}$ of $\Pi$, fixing any variable assignment of $\bar{x}$ with elements of $A$, we obtain that the truth of $\psi\vee\neg\phi_{1}\vee\cdots\vee\neg\phi_{m}$ under the instantiation with $R_{1},\ldots,R_{k}$ implies its truth under the instantiation with $R^{\prime}_{1},\ldots,R^{\prime}_{k}$: On the one hand, should any $\neg\phi_{i}$ be true, this is immediate (both for EDB and IDB atoms). On the other hand, in case all $\neg\phi_{i}$ are false, then so is the one $\neg\phi_{j}$ where $\phi_{j}=\phi_{j}(\bar{y})$ is the rule’s “frontier guard”, i.e., it consists of an EDB atom where $\bar{y}$ contains at least all the variables of $\bar{z}$ from $\psi(\bar{z})=\psi$. Truth of $\phi_{j}(\bar{y})$ under the chosen assignment implies that the tuple assigned to $\bar{y}$ is guarded and therefore also the tuple assigned to $\bar{z}$ must be guarded. Yet, then, truth of $\psi(\bar{z})$ under the two IDB instantiations coincides, hence $\psi(\bar{z})$ must hold.
However, not every problem in GSO which is in Datalog can be expressed by a frontier-guarded Datalog program. To prove this we need the following definition (in contrast to Definition 4.1, we work with a single parameter, since this is sufficient for our purposes).
@@
Let $\mathfrak{B}$ be an $\omega$-categorical structure with finite relational signature $\tau$ and $k\in{\mathbb{N}}$. Let $s$ be the maximal arity of $\tau$. Then the canonical frontier-guarded Datalog program of width $k$ for $\mathfrak{B}$ is the subset of the canonical Datalog program of width $(k,k)$ for $\mathfrak{B}$ which contains all rules that are frontier-guarded.
If $\Pi$ is the canonical frontier-guarded Datalog program of width $k$ for $\mathfrak{B}$ then it follows from Proposition 5.5 below that $\llbracket\Pi^{\prime}\rrbracket\subseteq\llbracket\Pi\rrbracket$ for every frontier-guarded Datalog program $\Pi^{\prime}$ of width $(k,k)$ which is sound for the complement of $\operatorname{CSP}(\mathfrak{B})$. We need to adapt the existential pebble game on $\mathfrak{A}$ and $\mathfrak{B}$ to the guarded setting as well. In our informal definition of the existential pebble game, we only change the rules for Spoiler: we additionally require that when Spoiler removes pebbles from $\mathfrak{A}$, then all the pebbles that remain on $\mathfrak{A}$ must be guarded, i.e., there must exist a tuple in a relation of $\mathfrak{A}$ such that these pebbles must be from the entries of this tuple. The resulting game will be called the existential guarded $k$-pebble game. Formally, we again work exclusively with the concept of a winning strategy $\mathcal{H}$} for Duplicator (similarly as in Definition 4.11). A partial homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$ is called guarded if there exists a tuple in a relation of $\mathfrak{A}$ whose entries contain $\operatorname{dom}(h)$.
@@
The following can be shown analogously to the proof of Theorem 4 in [14]. We omit the proof because all the ideas are present in the proof of a more interesting theorem that we present in full detail in Section 5.5 (Theorem 5.27).
@@
  • The canonical frontier-guarded Datalog program of width $k$ for $\mathfrak{B}$ does not derive $\sf goal$} on $\mathfrak{A}$.
@@
Recall from Example 1.1 that $\operatorname{CSP}({\mathbb{Q}};<)$ is expressible in MSO; it is easy to see that its complement can be defined by a Datalog program (of width (2,3)).
@@
Proof 5.7.
Let $k\in{\mathbb{N}}$. By Proposition 5.5, it suffices to show that there exists a structure $\mathfrak{A}$ which has no homomorphism to $({\mathbb{Q}};<)$ but Duplicator has a winning strategy for the existential guarded $k$ pebble game on $\mathfrak{A},({\mathbb{Q}};<)$. Let $\mathfrak{A}$ be a directed cycle of length at least $k+1$. Let $\mathcal{H}$} be the set of all partial homomorphisms $h$ from $\mathfrak{A}$ to $({\mathbb{Q}};<)$ with a domain of size at most $k$. We claim that $\mathcal{H}$ is a winning strategy for the existential guarded $(\ell,k)$-pebble game. Clearly, $\mathcal{H}$} is closed under restrictions. Now, let $h$ be a partial homomorphism whose domain is contained in a tuple from a relation of $\mathfrak{A}$, that is, contained in $\{a_{1},a_{2}\}$ for two subsequent elements of $A$ on the cycle. Let $a_{3},\dots,a_{k}$ be elements of $A$. Since $|A|\geq k+1$ there must be some $b\in A\setminus\{a_{1},\dots,a_{k}\}$. Let $\prec$ be the order on $a_{1},\dots,a_{k}$ in which these elements are traversed on the cycle, starting with $b$. Note that $\prec$ is a linear extension of $^{\mathfrak{A}}$. We may then extend $h$ to a partial homomorphism $h^{\prime}$ from $\mathfrak{A}$ to $({\mathbb{Q}};<)$ such that $h(a_{i})<h(a_{j})$ if $a_{i}\prec a_{j}$, for all $i,j\in\{1,\dots,k\}$. Note that $h^{\prime}\in{\mathcal{H}}$; this finishes the proof.
@@ 5.2 Flag and Check
A general mechanism, called flag and check, to turn a query language into a potentially more powerful query language has been introduced by Bourhis, Krötzsch, and Rudolph [21], generalising earlier work of Rudolph and Krötzsch [43]. The idea is best described for some fundamental constraint satisfaction problems. It is easy to see that there is no $k$ such that $\operatorname{CSP}(\{0,1\};\neq)$ (i.e., graph 2-colorability) is defined by a Datalog program of width $(1,k)$. However, to decide whether a graph is 2-colorable, it suffices to show that there is no vertex $v$ which can reach $v$ via a path of odd length. Note that such a ‘check’ can be performed by a Datalog program of width $(1,2)$ once a ‘flag’ has been put on $v$. Similarly, there is no $k\in{\mathbb{N}}$ such that $\operatorname{CSP}({\mathbb{Q}};<)$ (i.e., digraph acyclicity) can be defined by a Datalog program of width $(1,k)$ [9]. However, to decide whether a graph contains a directed cycle, it suffices to find a vertex $v$ which can reach $v$ by a directed path with at least one edge. Again, after $v$ has been found and ‘flagged’ such a computation can be performed by a Datalog program of width $(1,2)$.
@@
Let $\tau$ be a finite set of relation and constant symbols. A flag-and-check $\tau$-program ($\tau$-FCP) of arity $m$ is a set of Datalog rules $\Pi$ with EDBs from $\tau$ and the IDBs $\{{\sf goal},P_{1},\dots,P_{k}\}$ where $\sf goal$} is a distinguished predicate of arity 0. Moreover, on top of constants from $\tau$, $\Pi$ may use extra distinguished constant symbols $\lambda_{1},\dots,\lambda_{m}$. If $\mathfrak{A}$ is a $\tau$-structure, then $\Pi^{\mathfrak{A}}$ is the set of all tuples $(a_{1},\dots,a_{m})\in A^{m}$ such that $\Pi$ derives $\sf goal$} on the $\tau\cup\{\lambda_{1},\dots,\lambda_{m}\}$-expansion $\mathfrak{A}^{\prime}$ of $\mathfrak{A}$ where $\lambda^{\mathfrak{A}^{\prime}}_{i}=a_{i}$ for all $i\in\{1,\dots,m\}$.
@@
We omit the reference to $\tau$ in the notation if the reference is clear from the context. It has been shown that every problem in Modeq can be expressed in MSO [43]. Similarly, one can show that GQ is contained in GSO. This result is implicit in [21], which is why we present the proof here for the convenience of the reader.
@@
An example of a query that can be expressed in GQ, but not in Modeq can be found in [21] (Example 2). An example of a Datalog query in MSO which cannot be expressed in Modeq has already been found in [43] (Example 6); in fact, the same query is not even expressible in GQ, as we will show in Section 5.5 (Example 5.31), so we present it in more detail in the following.
@@
$\forall z\exists U_{2},U_{3}\forall u,v\big((C(u,z)$ $\Rightarrow U_{3}(u))$
@@
5.3 The Expressive Power of Modeqs
The Modeq formalism generalises natural concepts that have been discovered independently in the constraint satisfaction literature; this section can be skipped by impatient readers, and consulted by those interested in the relationship to constraint satisfaction.
Certain flag-and-check programs have been studied in the context of the complexity of constraint satisfaction under the name peek arc consistency (PAC) [11], extending the famous (hyper-) arc consistency procedure (AC; see, e.g., [10, Section 8.4]). In the following, $\tau$ denotes a finite relational signature.
Definition 5.12.
For example, the PAC procedure for $\operatorname{CSP}(\{0,1\};\neq)$ derives $\sf goal$} on a finite structure $\mathfrak{A}$ if and only if $\mathfrak{A}$ has no homomomorphism to $(\{0,1\};\neq)$; similarly, the PAC procedure solves $\operatorname{CSP}({\mathbb{Q}};<)$. We prove that the PAC procedure can be expressed as a Modeq.
Lemma 5.13.
Proof 5.14.
By the theorem of Ryll-Nardzewski, the automorphism group of $\mathfrak{B}$ has finitely many orbits $O_{1},\dots,O_{n}$; choose representatives $b_{1},\dots,b_{n}$ from these orbits. For every $i\leq n$, let $\Pi_{i}$ be the Datalog program for the hyperarc consistency procedure for $\operatorname{CSP}(\mathfrak{B},\{b_{i}\})$ (which has width $(1,p)$; see Section 8.4 in [10]). Let $C_{i}$ be the unary IDB of $\Pi_{i}$ for the relation $\{b_{i}\}$, and let $\Pi^{\prime}_{i}$ be the Datalog program obtained from $\Pi_{i}$ by adding the rule $C_{i}(\lambda_{1})\;{:}{-}$ (without precondition). Since expressibility by Datalog programs of width $(1,p)$ is closed under intersection (Lemma 4.3), there is a Datalog program $\Pi$ of width $(1,p)$ expressing $\Pi^{\prime}_{1}\wedge\cdots\wedge\Pi^{\prime}_{n}$. Also note that $\Theta:=\exists x.\Pi$ holds on a $\tau$-structure $\mathfrak{A}$ if and only if there exists an element $a\in A$ such that for every $i\leq n$ the program $\Pi^{\prime}_{i}$ derives $\sf goal$} on $(\mathfrak{A},\{a\})$, which by definition is the case if and only if the PAC procedure derives $\sf goal$} on $\mathfrak{A}$. This shows the statement.
The following example demonstrates that the expressivity of Modeq extends well beyond that of PAC.
Example 5.15.
The following is based on an example from [25] of a structure such that the complement of the CSP of this structure cannot be solved by PAC, but can be solved by singleton arc consistency (SAC). [^7] We do not need the notion of singleton arc consistency in this article and therefore refrain from giving a full definition and refer the interested reader to [25]. Below we will show that the mentioned class is even in Modeq. Let $\mathfrak{B}$ be the structure with domain $\{-1,0,1,2\}$ and relations
$R^{\mathfrak{B}}$ $\coloneqq\{-1,0,1,2\}^{2}\setminus\{(-1,-1)\}$
$S^{\mathfrak{B}}$ $\coloneqq\{(0,1),(1,2),(2,0),(-1,-1)\}$
$R^{\mathfrak{A}}$ $\coloneqq\{(a,a^{\prime})\}$
Then there is no homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$, but the PAC procedure for $\operatorname{CSP}(\mathfrak{B})$ does not derive $\sf goal$} on $\mathfrak{A}$ [25]. However, the problem can be defined by the Modeq $\exists x_{1},x_{2}.\,\Pi$ where $\Pi$ is the following monadic flag-and-check program of arity two.
$U_{0}(\lambda_{1})$ $\;{:}{-}$ $U_{0}(\lambda_{2})$ $\;{:}{-}$
$U_{1}(z)$ $\;{:}{-}\;U_{0}(y),S(y,z)$ $U_{2}(y)$ $\;{:}{-}\;U_{0}(z),S(y,z)$
$U_{2}(z)$ $\;{:}{-}\;U_{1}(y),S(y,z)$ $U_{0}(y)$ $\;{:}{-}\;U_{1}(z),S(y,z)$
$U_{0}(z)$ $\;{:}{-}\;U_{2}(y),S(y,z)$ $U_{1}(y)$ $\;{:}{-}\;U_{2}(z),S(y,z)$
${\sf goal}$ $\;{:}{-}\;U_{0}(y),U_{1}(y),R(y,z),U_{0}(z),U_{1}(z)$
Claim. $\exists x_{1},x_{2}.\,\Pi$ evaluates to true on a finite $\{R,S\}$-structure $\mathfrak{A}$ if and only if there is no homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$. First suppose that there are $a_{1},a_{2}\in A$ such that $\Pi$ derives $\sf goal$} on the $\{\lambda_{1},\lambda_{2}\}$-expansion $\mathfrak{A}^{\prime}$ of $\mathfrak{A}$ where $\lambda_{i}^{\mathfrak{A}_{1}}=a_{i}$ for $i\in\{1,2\}$. This means that there are $b_{1},b_{2}\in A$ such that $(b_{1},b_{2})\in R^{\mathfrak{A}}$, and there are paths of net length 0 and 1 from $a_{1}$ to $b_{1}$, and paths of net length 0 and 1 from $a_{2}$ to $b_{2}$. In general, the existence of paths of net length 0 and $1$ between two vertices implies that both have to be mapped to 1-1. Therefore, any homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$ must map both $b_{1}$ and $b_{2}$ to 1-1, which is impossible since $(b_{1},b_{2})\in R^{\mathfrak{A}}$.
Now suppose that for any $a_{1},a_{2}\in A$ the program $\Pi$ does not derive $\sf goal$} on the structure $\mathfrak{A}^{\prime}$ defined above. That is, for every $(b_{1},b_{2})\in R^{\mathfrak{A}}$ there exists $i\in\{1,2\}$ such that the for all $u\in A$ that are connected to $b_{i}$ in the graph defined by $S^{\mathfrak{A}}$ there is no path of net length 0 and no path of net length $1$ to $b_{i}$. Hence, there must exist a path of net length $2$ from $u$ to $b_{i}$. We then define $h(u):=b_{i}+2\mod 3$. For all other $u\in A$, we define h(u):=1h(u):=-1. It is straightforward to verify that $h$ is a homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$. $\triangle$